# Weyl law for 1-cycles

### v1

@@ 2.2. Almgren-Pitts Min-Max Theory
Let $(M,g)$ be an $n$-dimensional triangulated Riemannian manifold with corners. Let $\mathcal{Z}_{k}(M,\partial M)$ be the space of mod $2$ flat $k$-cycles on $M$ relative to $\partial M$ and let $\mathcal{I}_{k}(M,\partial M)$ the corresponding space of flat $k$-chains, $0\leq k\leq n$. We also consider the spaces $\mathcal{Z}_{k}(M)$ and $\mathcal{I}_{k}(M)$ of mod $2$ absolute flat cycles and chains on $M$ respectively as defined in [6] and in [5][Section 4.2.26]. The mass functional on the previous spaces with respect to the metric $g$ is denoted by $\operatorname{\mathbf{M}}$. Notice that for a relative chain $\tau\in\mathcal{I}_{k}(M,\partial M)$,
@@
where given an absolute $k$-chain $\tau^{\prime}$, $[\tau^{\prime}]$ denotes its class as a relative $k$-chain in $(M,\partial M)$. The flat norm associated to $\operatorname{\mathbf{M}}$ is denoted by $\mathcal{F}$, which for absolute $k$-chains is defined as
@@
We denote $I=[0,1]$ and $I^{d}=[0,1]^{d}$ for every $d\in\mathbb{N}$. In addition, given $q\in\mathbb{N}$ we denote $I(q)$ the complex obtained by dividing $I$ into $q$ equal parts, and $I^{d}(q)=I(q)\times...\times I(q)$ ($d$ times) with the product structure. We say that $I^{d}(q)$ is obtained from $I^{d}$ by performing a $q$-refinement.
@@
The convention in the previous two definitions is different from that in [18], [17] and [12].
@@
Let $X$ be a cubical complex, let $F:X_{0}\to\mathcal{Z}_{k}(M,\partial M)$ ($\mathcal{I}_{k}(M,\partial M)$, $\mathcal{Z}_{k}(M)$ or $\mathcal{I}_{k}(M)$ respectively) be a discrete family and let $q$ be an odd natural number. We define the $q$-refinement of $F$ to be the discrete family $R^{q}F:X(q)_{0}\to\mathcal{I}_{k}(M)$ given by $R^{q}F(x)=F(v)$ where $v$ is the point of $X_{0}$ closest to $x\in X(q)_{0}$.
@@
We denote by $\lambda_{k}\in H^{n-k}(\mathcal{Z}_{k}(M,\partial M;\mathbb{Z}_{2});\mathbb{Z}_{2})$ the fundamental cohomology class of $\mathcal{Z}_{k}(M,\partial M;\mathbb{Z}_{2})$ (see [11] and [16][Section 3]). It is defined by the property that a continuous family $F:X\to\mathcal{Z}_{k}(M,\partial M;\mathbb{Z}_{2})$ is a sweepout of $M$ (i.e. the gluing homomorphism described in [11] is nontrivial) if and only if $F^{*}(\lambda_{k})\neq 0$.
@@
Let $X$ be a cubical complex. We say that a continuous map $F:X\to\mathcal{Z}_{k}(M,\partial M)$ is a $p$-sweepout of $(M,\partial M)$ by $k$-cycles if $F^{*}(\lambda_{k}^{p})\neq 0$. We denote $\mathcal{P}_{p}^{k}=\mathcal{P}_{p}^{k}(M,\partial M)$ the collection of $p$-sweepouts on $(M,\partial M)$ by $k$-cycles with no concentration of mass.
@@
Notice that in the definition of $p$-sweepout and of $p$-widths, $X$ can be any cubical complex; this means any subcomplex of the $q$-refinement of $[0,1]^{d}$ for $q$ and $d$ arbitrary. Nevertheless, as $\lambda_{k}^{p}\in H^{p(n-k)}(\mathcal{Z}_{k}(M))$ is an $(n-k)p$-cohomology class, we can expect the simplicial complexes of dimension $p(n-k)$ to be enough to detect $\lambda_{k}^{p}$. As any $m$-dimensional polyhedron embeds in $\mathbb{R}^{2m+1}$, we should expect to obtain the same definition for the $p$-widths if we restrict $d=2(n-k)p+1$. This is indeed true, as the following proposition shows.
@@
Denote $\overline{\mathcal{P}}_{p}^{k}$ the set of $F\in\mathcal{P}_{p}^{k}$ such that the domain $X$ of $F$ is a $p(n-k)$-dimensional subcomplex of $I^{2p(n-k)+1}(q)$ for some $q\in\mathbb{N}$. Let
@@
As $\overline{\mathcal{P}}_{p}^{k}\subseteq\mathcal{P}_{p}^{k}$, it is clear that $\omega_{p}^{k}(M,g)\leq\overline{\omega}_{p}^{k}(M,g)$. To show the other inequality, let $F:X\to\mathcal{Z}_{k}(M,\partial M)$ be an element of $\mathcal{P}_{p}^{k}$. As $F^{*}(\lambda_{k}^{p})\neq 0$, there exists some $p(n-k)$-cycle $z$ in $X$ such that $\langle\lambda_{k}^{p},F_{*}([z])\rangle=0$, where $\langle$,$\rangle$ , ⟩ denotes the pairing between cohomology and homology in $\mathcal{Z}_{k}(M,\partial M)$. By the equivalence between singular homology and simplicial homology, we can assume that $z$ is a polyhedral cycle in $X$. Denote $Z$ its underlying cubical subcomplex. As $\dim(Z)=p(n-k)$, by the PL embedding theorem we can see $Z$ embedded in $\mathbb{R}^{2p(n-k)+1}$ (as some piecewise linear polyhedron whose cells are arbitrary simplices, not necessarily equilateral). Notice that $F:Z\to\mathcal{Z}_{k}(M,\partial M)$ is a $p$-sweepout. We would like to replace the domain $Z$ by some polyhedral cycle $Z^{\prime}$ supported in a cubical grid of $\mathbb{R}^{2p(n-k)+1}$ so that we can see $Z^{\prime}$ as a subcomplex of $I^{2p(n-k)+1}(q)$ for some $q\in\mathbb{N}$. Let $B$ be some tubular neighborhood of $Z$ in $\mathbb{R}^{2p(n-k)+1}$ and $R:B\to Z$ the corresponding retraction. By the Federer-Fleming Deformation Theorem, we can find a polyhedral $k$-cycle $Z^{\prime}$ and a $(k+1)$-chain $T$ both supported in $B$ such that $\partial T=Z^{\prime}-Z$. Then $[Z^{\prime}]=[Z]$ in $B$ therefore $R_{*}([Z^{\prime}])=R_{*}([Z])=[Z]$ in $Z$ which implies that the map $F^{\prime}:Z^{\prime}\to\mathcal{Z}_{k}(M)$ given by $F^{\prime}(x)=F(R(x))$ is a $p$-sweepout because
@@
As a consequence of Lemma 2.18, in the rest of the paper a cubical complex will denote a $p$-dimensional subcomplex of $I^{2p+1}(q)$ for some $p,q\in\mathbb{N}$.
@@ 3.1. Definitions
Throughout this section, $(M,g)$ is a fixed $n$-dimensional triangulated Riemannian manifold with corners. We fix $G=\mathbb{Z}$ or $\mathbb{Z}_{p}$ for some prime number $p\in\mathbb{N}$ and we denote $\mathcal{Z}_{k}(M,\partial M)$ the space of flat $k$-chains with coefficients in $G$ (in the rest of the paper we take $G=\mathbb{Z}_{2}$, but the results of this section hold for more general $G$). We will introduce the notion of $\delta$-localized family which was first defined in [12][Section 2]. Given an arbitrary continuous family $F:X\to\mathcal{Z}_{k}(M,\partial M)$, there might be arbitrarily close points $x,y\in M$ such that $\mathcal{F}(F(x)-F(y))$ is very small (because of continuity) but $\operatorname{spt}(F(x)-F(y))$ is very large; in fact that support could be almost dense in $M$ (and its mass might be very large too). The property of $\delta$-localization prevents from this to happen, as it requires that if $x$ and $y$ are close then the difference $F(x)-F(y)$ is supported in a collection of small balls. We make this more precise in the definitions below.
@@
Remark 3.3.
When $\partial M\neq\emptyset$, by $\operatorname{spt}(F(x)-F(y))$ we mean $\operatorname{spt}(\overline{F}(x)-\overline{F}(y))$ where $\overline{F}(x),\overline{F}(y)\in\mathcal{I}_{k}(X)$, $[\overline{F}(x)]=F(x)$, $[\overline{F}(y)]=F(y)$ and $F(x)\llcorner\partial M=F(y)\llcorner\partial M=0$ (i.e. $\overline{F}(x)$ and $\overline{F}(y)$ are absolute $k$-chains which represent $F(x)$ and $F(y)$ and have minimal mass, being $\operatorname{\mathbf{M}}(F(x))=\operatorname{\mathbf{M}}(\overline{F}(x))$ and $\operatorname{\mathbf{M}}(F(y))=\operatorname{\mathbf{M}}(\overline{F}(y))$ according to (1)).
@@
We will often need to construct $\delta^{\prime}$-admissible (respectively $(N^{\prime},\delta^{\prime})$-admissible) families from $\delta$-admissible (respectively $(N,\delta)$-admissible) ones. When carrying out such constructions, we normally first need to deal with families of balls which are not pairwise disjoint. The following lemma tells us that we can always get an accurate $\delta$-admissible (resp. $(N,\delta)$-admissible) family out of that.
@@
This is a consequence of [12][Lemma 2.4] and [12][Lemma 2.5]. ∎
@@
By induction in $p$. For $p=1$, any $(N,\delta)$-localized family is automatically monotonously localized so the result holds for $c(1)=1$. Assume that the result holds for $p-1$. Let $F:X^{p}\to\mathcal{I}_{k}(M)$ be $(N,\delta)$-localized and for each cell $C$ of $X$ denote $\{U_{i}^{C}\}_{i\in I_{C}}$ the $(N,\delta)$-admissible family associated to $C$. Then by inductive hypothesis $F_{p-1}=F|_{X_{p-1}}$ is monotonously $(c(p-1)N,c(p-1)\delta)$-localized. Given $E$ cell of $X$ with $\dim(E)\leq p-1$, denote $\{\tilde{U}^{E}_{i}\}_{i\in\tilde{I}_{E}}$ its corresponding $(c(p-1)N,c(p-1)\delta)$-admissible family. Given $C\in\operatorname{Faces}_{p}(X)$, let $\{\tilde{U}^{C}_{i}\}_{i\in\tilde{I}_{C}}$ be an $(c(p)N,c(p)\delta)$ admissible family which covers
@@
By Lemma 3.5, this can be done provided $c(p)\geq 3(1+2pc(p-1))$, as the total number of balls in the previous union is $N(1+2pc(p-1))$ (being $2p$ the number of $(p-1)$-faces of $C$) and the sum of their radius is $(1+2pc(p-1))\delta$. Thus setting $c(p)=3(1+2pc(p-1))$ we can see that $F$ is $(c(p)N,c(p)\delta)$-monotonously localized in the above defined families $\{\tilde{U}_{i}^{C}\}_{i\in\tilde{I}_{C}}$ for each $C\in\operatorname{Faces}(X)$. ∎
For the Parametric Isoperimetric Inequality, the family of 00-cycles we consider is supported in a triangulated Riemannian manifold with corners $(M,g)$, which allows us to proceed by induction in the skeleta of $M$. Therefore, we will need a notion of $\delta$-localized families supported in a PL subcomplex $P$ of $M$.
Definition 3.8.
Let $P\subseteq M$ be a subcomplex. A $\delta$-admissible family in $P$ is a disjoint collection of metric balls $\{B_{i}\}_{i\in I}$ in $P$ such that
$\sum_{i\in I}\operatorname{rad}(B_{i})<\delta.$
The definitions of $\delta$-localized and $(N,\delta)$-localized family $F:X_{j}\to\mathcal{I}_{k}(P)$ (respectively $\mathcal{I}_{k}(P,P\cap\partial M)$, $\mathcal{Z}_{k}(P)$ and $\mathcal{Z}_{k}(P,P\cap\partial M)$) are identical to those in Definition 3.2.
Given an subcomplex $P\subseteq M$, a $\delta$-localized family $G:X\to\mathcal{I}_{k}(P)$ and a top cell $Q$ of $P$, $G$ induces a continuous family $\overline{G}_{Q}:X\to\mathcal{I}_{k}(Q,\partial Q)$ given by $\overline{G}_{Q}(x)=[G(x)\llcorner Q]$. We may wonder whether we can represent $\overline{G}_{Q}$ by a continuous $\delta$-localized family of absolute chains in $Q$. The first attempt is to consider the family $x\mapsto G(x)\llcorner Q$ but this one may not be continuous. However, it is possible to construct such family by only adding $G(x)\llcorner Q$ certain $k$-chains supported in $\partial Q$ as the next lemma shows.
@@
Let $(M,g)$ be a triangulated Riemannian manifold with corners, let $P\subseteq M$ be a subcomplex and let $Q$ be a top dimensional cell of $P$. Let $G:X\to\mathcal{I}_{k}(P)$ be a continuous $\delta$-localized family of absolute $k$-chains, $k\geq 1$. Then there exists a continuous $G_{Q}:X\to\mathcal{I}_{k}(Q)$ such that
@@
    Given a cell $C$ of $X$ and $x,y\in C$, $G_{Q}(x)-G_{Q}(y)$ is supported in $\{\tilde{U}_{i}^{C}\}_{i\in I_{C}}$ where $\tilde{U}_{i}^{C}=U_{i}^{C}\cap Q$. By applying Lemma 3.5 we can see that $G_{Q}:X\to\mathcal{I}_{k}(Q)$ is $3\delta$-localized.
The same result holds taking $P=M$ and taking $Q$ to be a closed embedded $n$-dimensional submanifold with corners of $M$.
@@
where $r_{i}^{E}:U_{i}^{E}\to\tilde{U}_{i}^{E}$ is a retraction.
@@
At the 00-skeleton we just define $G_{Q}(x)=G(x)\llcorner Q$. Suppose $G_{Q}$ was defined in the $j$-skeleton and let $E$ be an $(j+1)$-cell of $X$. Pick a vertex $v$ of $E$. As $G|_{E}$ is localized in $\{U_{i}^{E}\}_{i\in I_{E}}$, there exists $G_{E}$ continuous and supported there such that
@@
Corollary 3.10.
If $G:X\to\mathcal{I}_{k}(M,\partial M)$, $k\geq 1$, is a continuous $\delta$-localized family then there exists a continuous $6\delta$-localized family of $k$-cycles $\overline{G}:X\to\mathcal{I}_{k}(M)$ such that $[\overline{G}(x)]=G(x)$.
Proof.
Let $\tilde{M}$ denote the double of $M$, obtained by gluing $M$ with itself by its boundary. Consider the $2\delta$-localized family $\tilde{G}$ on $\tilde{M}$ obtained by gluing $G$ with itself at $\partial M$ and apply the previous lemma to $Q=M\subseteq\tilde{M}$. Set $\overline{G}=G_{Q}$. ∎
Corollary 3.11.
Lemma 3.9 still holds if we consider families of relative chains $F:X\to\mathcal{I}_{k}(P,P\cap\partial M)$.

### v2

@@ 2.2. Almgren-Pitts Min-Max Theory
Let $(M,g)$ be an $n$-dimensional piecewise smooth Riemannian manifold and let $\mathbb{G}=\mathbb{Z}$ or $\mathbb{Z}_{p}$ for some prime $p$. Let $\mathcal{Z}_{k}(M,\partial M)$ be the space of flat $k$-cycles with coefficients in $\mathbb{G}$ on $M$ relative to $\partial M$ and let $\mathcal{I}_{k}(M,\partial M)$ the corresponding space of flat $k$-chains, $0\leq k\leq n$. We also consider the spaces $\mathcal{Z}_{k}(M)$ and $\mathcal{I}_{k}(M)$ of absolute flat cycles and chains on $M$ with coefficients in $\mathbb{G}$ as defined in [Fleming66] and in [FedererGMT][Section 4.2.26]. The mass functional on the previous spaces with respect to the metric $g$ is denoted by $\operatorname{\mathbf{M}}$. Notice that for a relative chain $\tau\in\mathcal{I}_{k}(M,\partial M)$,
@@
where given an absolute $k$-chain $\tau^{\prime}$, $\tau^{\prime}]$ denotes its class as a relative $k$-chain in $(M,\partial M)$. In addition, $\operatorname{spt}(\tau)=\operatorname{spt}(\tau^{\prime}-\tau^{\prime}\llcorner\partial M)$ for $\tau^{\prime}\in\mathcal{I}_{k}(M)$ such that $\tau^{\prime}]=\tau$ (the definition is independent of the representative $\tau^{\prime}$). The flat norm associated to $\operatorname{\mathbf{M}}$ is denoted by $\mathcal{F}$, which for absolute $k$-chains is defined as
@@
We denote $I=[0,1]$ and $I^{d}=[0,1]^{d}$ for each $d\in\mathbb{N}$. In addition, given $q\in\mathbb{N}$ we denote $I(q)$ the complex obtained by dividing $I$ into $q$ equal parts, and $I^{d}(q)=I(q)\times...\times I(q)$ ($d$ times) with the product structure. We say that $I^{d}(q)$ is obtained from $I^{d}$ by performing a $q$-refinement.
@@
The convention in the previous two definitions is different from that in [MN], [Willmore] and [GL22].
@@
Let $X$ be a cubical complex, let $F:X_{0}\to\mathcal{Z}_{k}(M,\partial M)$ ($\mathcal{I}_{k}(M,\partial M)$, $\mathcal{Z}_{k}(M)$ or $\mathcal{I}_{k}(M)$ respectively) be a discrete family and let $q$ be an odd natural number. We define the $q$-refinement of $F$ to be the discrete family $R^{q}F:X(q)_{0}\to\mathcal{Z}_{k}(M,\partial M)$ given by $R^{q}F(x)=F(v)$ where $v$ is the point of $X_{0}$ closest to $x\in X(q)_{0}$.
@@
For $\mathbb{G}=\mathbb{Z}_{2}$, we denote by $\lambda_{k}\in H^{n-k}(\mathcal{Z}_{k}(M,\partial M;\mathbb{Z}_{2});\mathbb{Z}_{2})$ the fundamental cohomology class of $\mathcal{Z}_{k}(M,\partial M;\mathbb{Z}_{2})$ (see [GuthMinMax] and [LS][Section 3]). It is defined by the property that a continuous family $F:X\to\mathcal{Z}_{k}(M,\partial M;\mathbb{Z}_{2})$ is a sweepout of $M$ (i.e. the gluing homomorphism described in [GuthMinMax] is nontrivial) if and only if $F^{*}(\lambda_{k})\neq 0$.
@@
Let $X$ be a cubical complex and $\mathbb{G}=\mathbb{Z}_{2}$. We say that a continuous map $F:X\to\mathcal{Z}_{k}(M,\partial M;\mathbb{Z}_{2})$ is a $p$-sweepout of $(M,\partial M)$ by $k$-cycles if $F^{*}(\lambda_{k}^{p})\neq 0$. We denote $\mathcal{P}_{p}^{k}=\mathcal{P}_{p}^{k}(M,\partial M)$ the collection of $p$-sweepouts on $(M,\partial M)$ by $k$-cycles with no concentration of mass.
@@
Notice that in the definition of $p$-sweepout and of $p$-widths, $X$ can be any cubical complex; this means any subcomplex of the $q$-refinement of $0,1]^{d}$ for $q$ and $d$ arbitrary. Nevertheless, as $\lambda_{k}^{p}\in H^{p(n-k)}(\mathcal{Z}_{k}(M;\mathbb{Z}_{2});\mathbb{Z}_{2})$ is a p(nk)p(n-k)-cohomology class, we can expect the simplicial complexes of dimension p(nk)p(n-k) to be enough to detect $\lambda_{k}^{p}$. As any $m$-dimensional polyhedron embeds in $\mathbb{R}^{2m+1}$, we should expect to obtain the same definition for the $p$-widths if we restrict d=2(nk)p+1d=2(n-k)p+1. This is indeed true, as the following proposition shows.
@@
Denote $\overline{\mathcal{P}}_{p}^{k}$ the set of $F\in\mathcal{P}_{p}^{k}$ such that the domain $X$ of $F$ is a p(nk)p(n-k)-dimensional subcomplex of $I^{2p(n-k)+1}(q)$ for some $q\in\mathbb{N}$. Let
@@
As $\overline{\mathcal{P}}_{p}^{k}\subseteq\mathcal{P}_{p}^{k}$, it is clear that $\omega_{p}^{k}(M,g)\leq\overline{\omega}_{p}^{k}(M,g)$. To show the other inequality, let $F:X\to\mathcal{Z}_{k}(M,\partial M)$ be an element of $\mathcal{P}_{p}^{k}$. As $F^{*}(\lambda_{k}^{p})\neq 0$, there exists some p(nk)p(n-k)-cycle $z$ in $X$ such that $\langle\lambda_{k}^{p},F_{*}([z])\rangle=0$, where $\langle,\rangle$ denotes the pairing between cohomology and homology in $\mathcal{Z}_{k}(M,\partial M)$. By the equivalence between singular homology and simplicial homology, we can assume that $z$ is a polyhedral cycle in $X$. Denote $Z$ its underlying cubical subcomplex. As $\dim(Z)=p(n-k)$, by the PL embedding theorem we can see $Z$ embedded in $\mathbb{R}^{2p(n-k)+1}$ (as some piecewise linear polyhedron whose cells are arbitrary simplices, not necessarily equilateral). Notice that $F:Z\to\mathcal{Z}_{k}(M,\partial M)$ is a $p$-sweepout. We would like to replace the domain $Z$ by some polyhedral cycle $Z^{\prime}$ supported in a cubical grid of $\mathbb{R}^{2p(n-k)+1}$ so that we can see $Z^{\prime}$ as a subcomplex of $I^{2p(n-k)+1}(q)$ for some $q\in\mathbb{N}$. Let $B$ be some tubular neighborhood of $Z$ in $\mathbb{R}^{2p(n-k)+1}$ and $R:B\to Z$ the corresponding retraction. By the Federer-Fleming Deformation Theorem, we can find a polyhedral $k$-cycle $Z^{\prime}$ and a $(k+1)$-chain $T$ both supported in $B$ such that $\partial T=Z^{\prime}-Z$. Then $Z^{\prime}]=[Z]$ in $B$ therefore $R_{*}([Z^{\prime}])=R_{*}([Z])=[Z]$ in $Z$ which implies that the map $F^{\prime}:Z^{\prime}\to\mathcal{Z}_{k}(M)$ given by $F^{\prime}(x)=F(R(x))$ is a $p$-sweepout because
@@
As a consequence of Lemma 2.25, in the rest of the paper a cubical complex will denote a $p$-dimensional subcomplex of $I^{2p+1}(q)$ for some $p,q\in\mathbb{N}$. This assumption will simplify some of the proofs, but they could still be carried out without it.
@@ 3.1. Definitions
Throughout this section, $(M,g)$ is a fixed $n$-dimensional smooth Riemannian manifold with boundary $\partial M=\Sigma$. We fix $\mathbb{G}=\mathbb{Z}$ or $\mathbb{Z}_{p}$ for some prime number $p$ and we denote $\mathcal{Z}_{k}(M,\partial M)$ the space of flat relative $k$-cycles (respectively relative chains, and absolute cycles and chains) with coefficients in $\mathbb{G}$ as introduced in the previous section (in the rest of the paper we take $\mathbb{G}=\mathbb{Z}_{2}$, but the results of this section hold for more general $\mathbb{G}$). In fact, all the arguments of this section also work by replacing the pair $(M,\partial M)$ by a pair $(P,P^{\prime})$ where $P$ is an $n$-dimensional Riemannian polyhedron and $P^{\prime}$ is a subpolyhedron of $P$. This will be used in Section 4.3 when proving the Parametric Isoperimetric Inequality for 0-cycles. We will introduce the notion of $\delta$-localized family which was first defined in [GL22][Section 2]. Given an arbitrary continuous family $F:X\to\mathcal{Z}_{k}(M,\partial M)$, there might be arbitrarily close points $x,y\in X$ such that $\mathcal{F}(F(x)-F(y))$ is very small (because of continuity) but $\operatorname{spt}(F(x)-F(y))$ is very large; in fact the support of F(x)F(y)F(x)-F(y) could be almost dense in $M$ (and its mass might be very large too). The property of $\delta$-localization prevents from this to happen, as it requires that if $x$ and $y$ are close then the difference F(x)F(y)F(x)-F(y) is supported in a collection of small balls. We make this more precise in the definitions below.
@@
Remark 3.2.
If $P$ is a Riemannian polyhedron, we just require the $U_{i}$ to be metric balls such that $\sum_{i\in I}r_{i}<\delta$.
@@
We will often need to construct $\delta^{\prime}$-admissible (respectively, $(N^{\prime},\delta^{\prime})$-admissible) families from $\delta$-admissible (respectively $(N,\delta)$-admissible) ones. When carrying out such constructions, we normally first need to deal with families of balls which are not pairwise disjoint. The following lemma tells us that we can always get an accurate $\delta$-admissible (resp. $(N,\delta)$-admissible) family out of that.
@@
This is a consequence of [GL22][Lemma 2.4] and [GL22][Lemma 2.5]. ∎
@@
By induction in $p$. For $p=1$, any $(N,\delta)$-localized family is automatically monotonously localized so the result holds for $c(1)=1$. Assume that the result holds for p1p-1. Let $F:X^{p}\to\mathcal{I}_{k}(M)$ be $(N,\delta)$-localized and for each cell $C$ of $X$ denote $\{U_{i}^{C}\}_{i\in I_{C}}$ the $(N,\delta)$-admissible family associated to $C$. Then by inductive hypothesis $F_{p-1}=F|_{X_{p-1}}$ is monotonously $(c(p-1)N,c(p-1)\delta)$-localized. Given $E$ cell of $X$ with $\dim(E)\leq p-1$, denote $\{\tilde{U}^{E}_{i}\}_{i\in\tilde{I}_{E}}$ its corresponding $(c(p-1)N,c(p-1)\delta)$-admissible family. Given $C\in\operatorname{Faces}_{p}(X)$, let $\{\tilde{U}^{C}_{i}\}_{i\in\tilde{I}_{C}}$ be an $(c(p)N,c(p)\delta)$ admissible family which covers
@@
By Lemma 3.5, this can be done provided $c(p)\geq 3(1+2pc(p-1))$, as the total number of balls in the previous union is N(1+2pc(p1))N(1+2pc(p-1)) (being $2p$ the number of (p1)(p-1)-faces of $C$) and the sum of their radius is at most $(1+2pc(p-1))\delta$. Thus setting c(p)=3(1+2pc(p1))c(p)=3(1+2pc(p-1)) we can see that $F$ is $(c(p)N,c(p)\delta)$-monotonously localized in the above defined families $\{\tilde{U}_{i}^{C}\}_{i\in\tilde{I}_{C}}$ for each $C\in\operatorname{Faces}(X)$. ∎
Given a Riemannian polyhedron $P$, a $\delta$-localized family $G:X\to\mathcal{I}_{k}(P)$ and a top cell $Q$ of $P$, $G$ induces a continuous family $\overline{G}_{Q}:X\to\mathcal{I}_{k}(Q,\partial Q)$ of relative $k$-chains given by $\overline{G}_{Q}(x)=[G(x)\llcorner Q]$. We may wonder whether we can represent $\overline{G}_{Q}$ by a continuous $\delta$-localized family of absolute chains in $Q$. The first attempt is to consider the family $x\mapsto G(x)\llcorner Q$ but this one may not be continuous. However, it is possible to construct such family by only adding $G(x)\llcorner Q$ certain $k$-chains supported in $\partial Q$ as the next lemma shows.
@@
Let $P$ be a Riemannian polyhedron and let $Q$ be a top dimensional cell of $P$. Let $G:X\to\mathcal{I}_{k}(P)$ be a continuous $\delta$-localized family of absolute $k$-chains, $k\geq 1$. Then there exists a continuous $G_{Q}:X\to\mathcal{I}_{k}(Q)$ such that
@@
    Given a cell $C$ of $X$ and $x,y\in C$, $G_{Q}(x)-G_{Q}(y)$ is supported in $\{\tilde{U}_{i}^{C}\}_{i\in I_{C}}$ where $\tilde{U}_{i}^{C}=U_{i}^{C}\cap Q$. Thus $G_{Q}$ is $\delta$-localized.
The same result holds taking $P=M$ and taking $Q$ to be a closed embedded $n$-dimensional piecewise smooth Riemannian submanifold of $M$.
@@
where $r_{i}^{E}:U_{i}^{E}\to\tilde{U}_{i}^{E}$ is a Lipschitz retraction mapping $U_{E}\cap(M\setminus Q)$ onto $U_{E}\cap\partial Q$.
@@
At the 0-skeleton we just define $G_{Q}(x)=G(x)\llcorner Q$. Suppose $G_{Q}$ was defined in the $j$-skeleton and let $E$ be an $(j+1)$-cell of $X$. Pick a vertex $v$ of $E$. As $G|_{E}$ is localized in $\{U_{i}^{E}\}_{i\in I_{E}}$, there exists $G_{E}$ continuous and supported there such that